OpenStudy (anonymous):

how do you find the second endpoint when given the midpoint and one of the endpoints? is there a formula or what?

OpenStudy (aravindg):

you see midpoint is got by taking arithmetic mean of end points

OpenStudy (aravindg):

so call the unknown endpoint as (x,y) if known endpoint is (a,b)

OpenStudy (aravindg):

midpoint (b,c)=(x+a/2,y+b/2)

OpenStudy (aravindg):

2b=x+a 2c=y+b

OpenStudy (perl):

hellooooooo!!!

OpenStudy (aravindg):

then you can find (x,y)

OpenStudy (aravindg):

gt it ?

OpenStudy (anonymous):

thank you @AravindG

OpenStudy (aravindg):

yw :)

OpenStudy (perl):